# Survey of Video Diffusion Models: Foundations, Implementations, and Applications

### v1

@@ 2.4.4 Agent-based diffusion models
VideoGPT(Yan et al., 2021) - VQVAE - - UCF-101, BAIR $64\times 64$
CogVideo(Hong et al., 2023a) DiT 2D VAE GPT-3 15.5B WebVid-5.4M $480\times 480$ 8 $\times$ RTX 6000 -
@@
MagicVideo(Zhou et al., 2022) 3D UNet VideoVAE CLIP - WebVid-10M, HD-VG-130M $1024\times 1024$ 1 $\times$ A100 -
Make-A-Video(Singer et al., 2022) UNet 2D VAE CLIP 9.7B WebVid-10M, HD-VG-100M $768\times 768$ 76 - -
LVDM(Rakhimov et al., 2023) 3D UNet 3D VAE CLIP 1.6B UCF-101, TaiChi $128\times 128$ 1024 8 $\times$ V100s 4.5 days
@@
MagViT-V2(Yu et al., 2024b) - 3D VAE MLM 307M UCF-101 $640\times 360$
VideoFusion(Zhou et al., 2023) UNet - CLIP 2B UCF101, Taichi-HD, SkyTimelapse $128\times 128$ 16 - -
VideoComposer(Jiang et al., 2023a) 3D UNet - CLIP - WebVid-10M $256\times 256$ 16 - -
@@
ModelScope(Wang et al., 2023c) 3D UNet VQGAN T5 $\sim$1.7B WebVid $256\times 256$ A100 -
HiGen(Qing et al., 2024) 3D UNet - CLIP - WebVid-10M $448\times 256$ 32 8 $\times$ A100s -
@@
VideoGen(Chen & Sun, 2023) UNet - CLIP - WebVid-10M $256\times 256$ 16 64 $\times$ A100s -
Animate-A-Story(He et al., 2023a) 3D UNet - CLIP - WebVid-10M $256\times 256$ 16 - -
AnimateDiff(Guo et al., 2023c) UNet 2D VAE - - Web Vid $1024\times 576$, -
AnimateDiff-V2 UNet - CLIP - - $1024\times 1024$, -
SimDA(Xing et al., 2024b) UNet - T5-XXL 1.1B WebVid-10M $256\times 256$ 1 $\times$ A100 -
@@
Snap Video(Menapace et al., 2024) DiT 3D VAE T5-XXL 3.90B UCF-101, MSRVTT $288\times 512$ 16 A100 -
@@
SiT(Ma et al., 2024a) DiT 2D VAE CLIP ViT-L 675M LAION-5B $512\times 512$
Open-Sora(Zheng et al., 2024c) STDiT 3D VAE T5-XXL 1.1B WebVid-10M, Panda-70M HD-VG-130M, MiraData Vript, Inter4K $1280\times 720$ 2-16s 16 8 $\times$ H100s 3.5k hrs
Open-Sora-Plan(Lab & etc., 2024) DiT WF-VAE mT5-XXL - Panda-70M $256\times 256$ 25-49 8×8$\times$8 × NPUs -
@@
W.A.L.T(Gupta et al., 2023) DiT - T5-XXL 3.00B UCF-101, MSRVTT $512\times 896$ 2-5s - 8×8$\times$8 × A6000s -
Phenaki(Villegas et al., 2022) C-ViViT 2D VAE T5-XXL 0.8B $64\times 64$ 11-15 - -
@@
Seine (Chen et al., 2024d) UNet VQGAN CLIP - WebVid-10M $320\times 512$ 16 - -
@@
Optis (Chen et al., 2024d) - - CLIP 1.3B WebVid-10M $320\times 512$ 5min 320 A100 -
@@
Show-1 UNet - T5-XXL - WebVid-10M $576\times 320$ 48 $\times$ A100s -
@@
VideoCrafter1(Chen et al., 2023a) U-ViT Video VAE CLIP - Laion-coco-600M, Webvid-10M $1024\times 576$
VideoCrafter2(Chen et al., 2024a) U-ViT Video VAE - - WebVid-10M 512 $\times$ 320 - - -
@@
Step-Video-T2V (Ma et al., 2025) DiT Video VAE CLIP, Step-LLM $\sim$30B - $768\times 768$ 204 H100 -
SkyReels-V1 (SkyReels-AI, 2025) DiT 3D VAE MLLM - - $544\times 960$ 97 - -
@@
Table 1: Comparison of modules and parameters in different diffusion generative models and their industry applications. New models can be found in video generation arena leaderboard and vbench leaderboard. (TAE refers to temporal autoencoder, and MLLM refers to multimodal large language model). Papers about related component and dataset: VideoVAE(Zhou et al., 2022), mT5-XXL(Xue, 2020), WebVid(Bain et al., 2021), UCF-101(Soomro et al., 2012), MSR-VTT(Xu et al., 2016), MiraData(Ju et al., 2024), Laion-coco 600M(LAION, 2023) Taichi-HD(Siarohin et al., 2019), SkyTimelapse(Xiong et al., 2018), Panda-70M(Chen et al., 2024c), FaceForensics(Rössler et al., 2018), Inter4K(Stergiou & Poppe, 2023), Vimeo25M(Huang et al., 2023b), MiraData(Zhao et al., 2023d), Vript(Yang et al., 2024a)
2.5 Architecture
As shown in Figure 4, most text-conditional visual generation models comprise three key modules: a variational autoencoder (VAE) compresses images and videos into a latent space. Second, a neural network—typically incorporating a U-Net or Transformer-based backbone—performs the denoising within this latent space. The denoising process can be enhanced using optical flow or human feedback to improve generation quality. A text encoder generates embeddings to guide the image or video generation process.
@@ 2.5.1 UNet
UNet has been further adapted to handle temporal information for video-based diffusion tasks. Extensions like AnimateDiff incorporate temporal attention layers within the UNet to capture dependencies across frames. (An et al., 2023) keep using 2D UNet and enable motion learning by shifting the feature channels along the temporal dimension. Other modifications include transitioning from classic 2D-UNet to spatial-temporal factorized 3D UNet architecture. VDM (Ho et al., 2022c) propose a spatial-temporal factorized 3D UNet by adding temporal blocks for video generation, as a natural extension of the standard image diffusion model. Similarly, (Ho et al., 2022a) builds on the similar video U-Net architecture to the cascaded image diffusion model Imagen Saharia et al. (2022), while (He et al., 2022), MagicVideo(Zhou et al., 2022), and ModelScope(Wang et al., 2023c)apply it to the latent space (Rombach et al., 2022), as shown in Table 1. MoVideo(Liang et al., 2025) further generates the depth and optical flow of the whole video by utilizing a 3D UNet architecture by adding extra-temporal modules, including temporal convolution and temporal attention layers after spatial convolution and spacial attention layers, improving frame consistency and visual quality for video generation.
@@ 2.5.2 Diffusion transformers
The Diffusion Transformer (DiT) replaces UNet’s traditional convolutional design with a Vision Transformer (ViT) framework, offering improved spatial attention capabilities suited for generative tasks across images and videos(Peebles & Xie, 2023). DiT achieves this through a 2D patchification process, where individual image frames are divided into patches that are then treated as tokens, allowing the model to capture detailed spatial dependencies within each frame. GenTrone(Chen et al., 2024b) adapted DiT from class to text conditioning and scaled GenTrone from approximately 900M to over 3B parameters demonstrating Transformer-based diffusion models can be used in the visual generative domain. From Table 1, models including CausVid(Yin et al., 2024), SiT(Ma et al., 2024a), and Sora also use DiT in video generation settings. However, for video generation, 2D tokenization alone is insufficient to capture the sequential dependencies critical for temporal coherence across frames. Approaches such as ViViT(Arnab et al., 2021) and STDiT(Zheng et al., 2024c) address this limitation by extending DiT’s patchification from 2D to 3D, introducing a 3D tokenization approach that simultaneously captures both spatial and temporal relationships. One representative work is CogVideoX (Hong et al., 2023a), which uses STDiT to model spatial-temporal evolution in video generation, capturing long-range dependencies for improved motion consistency and semantic fidelity. In this structure, consecutive frames are processed as spatiotemporal patches, enabling the model to understand dynamic changes across frames. Building on these approaches, frameworks like VDT(Lu et al., 2024a) incorporate both causal attention, which restricts attention to previous frames, and sparse causal attention, which further limits the focus to a subset of recent frames. VersVideo(Xiang et al., 2023) incorporated multi-excitation paths for spatial-temporal convolutions with dimension pooling across different axes and multi-expert spatial-temporal attention blocks to further boost the model’s spatial-temporal performance instead of simply extending 2D operations with temporal operations to significantly escalate training and inference costs.Text2Performer(Jiang et al., 2023b)proposes ae continuous VQ-diffuser to directly output the continuous pose embeddings for better motion modeling. These mechanisms enhance model’s effectiveness for video-based diffusion tasks.
@@ 2.5.3 VAE for latent space compression
Classical VAEs for image and video compression include standard VAEs(Kingma, 2013), quantized versions like VQVAE(Esser et al., 2021), VQGAN(Van Den Oord et al., 2017), and their GAN-enhanced variants, which improve reconstruction quality for higher compression. For example, VideoGPT(Yan et al., 2021) use a 3D-VQVAE in the latent space for video generation. Seine(Chen et al., 2024d) and MAGVIT(Yu et al., 2023a) integrate 3D-VQGAN with discriminators for latent space encoding to achieve better visual quality. The VAE framework not only compresses effectively but also enable the training of multiple generative tasks across downstream tasks.
@@
By integrating VAE in model building, some image generation models(Ho et al., 2022a; Li et al., 2024b; Podell et al., 2023; Bao et al., 2023b; Peebles & Xie, 2023; Lu et al., 2024c; Ma et al., 2024a; Chen et al., 2023c; Li et al., 2024j; Team, 2024b; Esser et al., 2024) also leverage latent space encoding/decoding, freezing VAE parameters during diffusion training and inference. Some diffusion models, including Make-A-Video(Singer et al., 2022), Imagen(Ho et al., 2022a), Show-1(Zhang et al., 2023a), directly learn pixel distributions generating videos, as shown in Table 1. However, video generation involves both spatial and temporal complexity, leading to higher computational costs. Besides, some representative diffusion video generation models, including Sora(Liu et al., 2024e), Phenaki(Villegas et al., 2022), VideoCrafter(Chen et al., 2023a), AnimateDiff(Guo et al., 2023c), and VideoPoet(Kondratyuk et al., 2023), start using VAE compression in video settings before training and inference in latent space. Many video generation models, including Latte(Ma et al., 2024c), Lavie(Huang et al., 2023b), MagicVideo(Zhou et al., 2022), Align-your-latent(Blattmann et al., 2023b), and AnimateDiff(Guo et al., 2023c), are devrived from Stable Diffusion’s image 2D VAEs, as training a full 3D latent space from scratch is challenging. However, temporal compression in 2D VAE-based video models relies on uniform frame sampling, which ignores motion information, leading to low FPS and less smooth video generation.
@@
Some approaches also start to utilize hybrid 2D-3D or fully 3D VAEs. For example, MagViT(Yu et al., 2023a), EasyAnimate(Xu et al., 2024c), Open-Sora(Zheng et al., 2024c; Lab & etc., 2024), and CogVideo-X(Yang et al., 2024j) employ hybrid 2D-3D VAEs, leveraging volumetric encoding for video generation. MAGViT(Yu et al., 2023a) adopting 3D VQGAN structures integrating both 3D and 2D downsampling, and MAGViT-V2(Yu et al., 2024b) utilizing a fully 3D VAE with 3D convolutional encoder and overlapping downsampling for enhanced spatial-temporal fidelity. HunyuanVideo(Kong et al., 2025), Kling, Mochi 1 (Team, 2024a), and SkyReels-V1(SkyReels-AI, 2025) continue leveraging 3D VAEs for compressed training in latent space in industry scenarios. Step-Video-T2V(Ma et al., 2025) and Magic-1-For-1(Yi et al., 2025) introduce novel Video VAE structures for efficient generation. However, training a video VAE independently without ensuring compatibility often leads to a latent space gap, preventing accurate projection into pixel space (Zhao et al., 2024b). CV-VAE (Zhao et al., 2024b) introduces a novel approach using latent space regularization to train a video VAE that extracts continuous latent for generative video models while maintaining compatibility with existing pretrained image and video models. Also, to trade off lower memory and computational cost with slightly lower reconstruction quality, the latest video generation model, Moive Gen(teamMeta, 2024) further uses interleaved 2D-1D convolutional encoders in its VAE. This approach balances efficiency and quality by reducing the redundancy of temporal information while maintaining spatial coherence.
2.5.4 Text encoder

### v2

@@ 2.4.4 Agent-based diffusion models
VideoGPT(Yan et al., 2021) - VQVAE - - UCF-101, BAIR $64\times 64$ - - - -
CogVideo(Hong et al., 2023a) DiT 2D VAE GPT-3 15.5B WebVid-5.4M $480\times 480$ - - 8 $\times$ RTX 6000 -
@@
MagicVideo(Zhou et al., 2022) 3D UNet VideoVAE CLIP - WebVid-10M, HD-VG-130M $1024\times 1024$ - - 1 $\times$ A100 -
Make-A-Video(Singer et al., 2022) UNet 2D VAE CLIP 9.7B WebVid-10M, HD-VG-100M $768\times 768$ - 76 - -
LVDM(Rakhimov et al., 2023) 3D UNet 3D VAE CLIP 1.6B UCF-101, TaiChi $128\times 128$ - 1024 8 $\times$ V100s 4.5 days
@@
MagViT-V2(Yu et al., 2024b) - 3D VAE MLM 307M UCF-101 $640\times 360$ - - - -
VideoFusion(Zhou et al., 2023) UNet - CLIP 2B UCF101, Taichi-HD, SkyTimelapse $128\times 128$ - 16 - -
VideoComposer(Jiang et al., 2023a) 3D UNet - CLIP - WebVid-10M $256\times 256$ - 16 - -
@@
ModelScope(Wang et al., 2023c) 3D UNet VQGAN T5 $\sim$1.7B WebVid $256\times 256$ - - A100 -
HiGen(Qing et al., 2024) 3D UNet - CLIP - WebVid-10M $448\times 256$ - 32 8 $\times$ A100s -
@@
VideoGen(Chen & Sun, 2023) UNet - CLIP - WebVid-10M $256\times 256$ - 16 64 $\times$ A100s -
Animate-A-Story(He et al., 2023a) 3D UNet - CLIP - WebVid-10M $256\times 256$ - 16 - -
AnimateDiff(Guo et al., 2023c) UNet 2D VAE - - Web Vid $1024\times 576$ - - -, -
AnimateDiff-V2 UNet - CLIP - - $1024\times 1024$ - - -, -
SimDA(Xing et al., 2024b) UNet - T5-XXL 1.1B WebVid-10M $256\times 256$ - - 1 $\times$ A100 -
@@
Snap Video(Menapace et al., 2024) DiT 3D VAE T5-XXL 3.90B UCF-101, MSRVTT $288\times 512$ - 16 A100 -
@@
SiT(Ma et al., 2024a) DiT 2D VAE CLIP ViT-L 675M LAION-5B $512\times 512$ - - - -
Open-Sora(Zheng et al., 2024c) STDiT 3D VAE T5-XXL 1.1B
WebVid-10M, Panda-70M
HD-VG-130M, MiraData
Vript, Inter4K
$1280\times 720$ 2-16s 16 8 $\times$ H100s 3.5k hrs
Open-Sora-Plan(Lab & etc., 2024) DiT WF-VAE mT5-XXL - Panda-70M $256\times 256$ - 25-49 $8\times$ NPUs -
@@
W.A.L.T(Gupta et al., 2023) DiT - T5-XXL 3.00B UCF-101, MSRVTT $512\times 896$ 2-5s - $8\times$ A6000s -
Phenaki(Villegas et al., 2022) C-ViViT 2D VAE T5-XXL 0.8B $64\times 64$ - 11-15 - -
@@
Seine (Chen et al., 2024e) UNet VQGAN CLIP - WebVid-10M $320\times 512$ - 16 - -
@@
Optis (Chen et al., 2024e) - - CLIP 1.3B WebVid-10M $320\times 512$ 5min 320 A100 -
@@
Show-1 UNet - T5-XXL - WebVid-10M $576\times 320$ - - 48 $\times$ A100s -
@@
VideoCrafter1(Chen et al., 2023a) U-ViT Video VAE CLIP - Laion-coco-600M, Webvid-10M $1024\times 576$ - - - -
VideoCrafter2(Chen et al., 2024b) U-ViT Video VAE - - WebVid-10M 512 $\times$ 320 - - -
@@
Step-Video-T2V (Ma et al., 2025) DiT Video VAE CLIP, Step-LLM $\sim$30B - $768\times 768$ - 204 H100 -
SkyReels-V1 (SkyReels-AI, 2025) DiT 3D VAE MLLM - - $544\times 960$ - 97 - -
@@
Table 1: Comparison of modules and parameters in different diffusion generative models and their industry applications. New models can be found in video generation arena leaderboard and vbench leaderboard. (TAE refers to temporal autoencoder, and MLLM refers to multimodal large language model). Papers about related component and dataset: VideoVAE(Zhou et al., 2022), mT5-XXL(Xue, 2020), WebVid(Bain et al., 2021), UCF-101(Soomro et al., 2012), MSR-VTT(Xu et al., 2016), MiraData(Ju et al., 2024), Laion-coco 600M(LAION, 2023) Taichi-HD(Siarohin et al., 2019), SkyTimelapse(Xiong et al., 2018), Panda-70M(Chen et al., 2024d), FaceForensics(Rössler et al., 2018), Inter4K(Stergiou & Poppe, 2023), Vimeo25M(Huang et al., 2023b), MiraData(Zhao et al., 2023d), Vript(Yang et al., 2024a).
2.5 Architectures
As shown in Figure 4, most text-conditional visual generation models comprise three key modules: a variational autoencoder (VAE) compresses images and videos into a latent space. Second, a neural network, typically incorporating a U-Net or Transformer-based backbone, performs the denoising within this latent space. The denoising process can be enhanced using optical flow or human feedback to improve generation quality. A text encoder generates embeddings to guide the image or video generation process. These are the main architectures in video generation tasks as shown in Figure 1.
@@ 2.5.1 UNet
UNet has been further adapted to handle temporal information for video-based diffusion tasks. Extensions like AnimateDiff incorporate temporal attention layers within the UNet to capture dependencies across frames. (An et al., 2023) keep using 2D UNet and enable motion learning by shifting the feature channels along the temporal dimension. Other modifications include transitioning from classic 2D-UNet to spatial-temporal factorized 3D UNet architecture. VDM (Ho et al., 2022c) propose a spatial-temporal factorized 3D UNet by adding temporal blocks for video generation, as a natural extension of the standard image diffusion model. Similarly, (Ho et al., 2022a) builds on the similar video U-Net architecture to the cascaded image diffusion model Imagen Saharia et al. (2022), while (He et al., 2022), MagicVideo(Zhou et al., 2022), and ModelScope(Wang et al., 2023c)apply it to the latent space (Rombach et al., 2022), as shown in Table 1. MoVideo(Liang et al., 2025a) further generates the depth and optical flow of the whole video by utilizing a 3D UNet architecture by adding extra-temporal modules, including temporal convolution and temporal attention layers after spatial convolution and spacial attention layers, improving frame consistency and visual quality for video generation.
To address the challenge of maintaining temporal consistency across video frames, modern video diffusion models incorporate temporal attention mechanisms within their UNet architectures. Temporal attention enables the model to capture dependencies between frames by computing attention weights across the temporal dimension. This mechanism is particularly crucial for video generation tasks where maintaining object consistency and smooth motion transitions is essential. The temporal attention operation can be formulated as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (13)
where $Q$, $K$, and $V$ represent query, key, and value tensors derived from different temporal positions, and $d_{k}$ is the dimension of the key vectors. This formulation allows the model to selectively attend to relevant temporal information, enhancing the quality of generated videos by ensuring coherent motion and consistent object appearances across frames.
One notable implementation is Align Your Latents(Blattmann et al., 2023b), which introduces a sophisticated temporal attention mechanism that operates in the latent space, significantly improving the temporal coherence of generated videos while maintaining high spatial resolution. As illustrated in Figure 5, this approach demonstrates how temporal attention can be effectively integrated into existing diffusion architectures to enhance video generation capabilities.
Figure 5: Classic temporal attention design from (Blattmann et al., 2023b), showing how temporal attention mechanisms improve video generation quality by maintaining temporal coherence across frames.
@@ 2.5.2 Diffusion transformers
The Diffusion Transformer (DiT) replaces UNet’s traditional convolutional design with a Vision Transformer (ViT) framework, offering improved spatial attention capabilities suited for generative tasks across images and videos (Peebles & Xie, 2023). DiT achieves this through a 2D patchification process, where individual image frames are divided into patches that are then treated as tokens, allowing the model to capture detailed spatial dependencies within each frame. GenTrone(Chen et al., 2024c) adapted DiT from class to text conditioning and scaled GenTrone from approximately 900M to over 3B parameters demonstrating Transformer-based diffusion models can be used in the visual generative domain. From Table 1, models including CausVid(Yin et al., 2024), SiT(Ma et al., 2024a), and Sora also use DiT in video generation settings. However, for video generation, 2D tokenization alone is insufficient to capture the sequential dependencies critical for temporal coherence across frames. Approaches such as ViViT(Arnab et al., 2021) and STDiT(Zheng et al., 2024c) address this limitation by extending DiT’s patchification from 2D to 3D, introducing a 3D tokenization approach that simultaneously captures both spatial and temporal relationships. One representative work is CogVideoX (Hong et al., 2023a), which uses STDiT to model spatial-temporal evolution in video generation, capturing long-range dependencies for improved motion consistency and semantic fidelity. In this structure, consecutive frames are processed as spatiotemporal patches, enabling the model to understand dynamic changes across frames. Building on these approaches, frameworks like VDT(Lu et al., 2024a) incorporate both causal attention, which restricts attention to previous frames, and sparse causal attention, which further limits the focus to a subset of recent frames. VersVideo(Xiang et al., 2023) incorporated multi-excitation paths for spatial-temporal convolutions with dimension pooling across different axes and multi-expert spatial-temporal attention blocks to further boost the model’s spatial-temporal performance instead of simply extending 2D operations with temporal operations to significantly escalate training and inference costs. Text2Performer(Jiang et al., 2023b)proposes ae continuous VQ-diffuser to directly output the continuous pose embeddings for better motion modeling. These mechanisms enhance model’s effectiveness for video-based diffusion tasks.
The transition from 2D to 3D patchification represents a fundamental architectural innovation for video generation. ViViT (Arnab et al., 2021) introduces a factorized encoder architecture that separates spatial and temporal modeling. As shown in Figure 6, this model first extracts patch tokens from each frame using spatiotemporal tokenization. A spatial transformer encoder then processes the tokens within each frame independently to model local structure and appearance.
The model passes the outputs of each frame to a temporal transformer that captures how frames relate over time. This two-step design learns both spatial and temporal patterns clearly and separately. For video generation, this setup helps control motion and appearance more precisely and keeps the frames consistent over time. Separating spatial and temporal steps also makes the model easier to train and understand.
Figure 6: Factorized encoder (Model 2) in ViViT (Arnab et al., 2021). The first transformer encoder models spatial interactions within each frame. The second transformer encoder captures temporal dependencies between frames. This factorized structure enables a late fusion of spatial and temporal information, which supports temporally consistent video generation.
@@ 2.5.3 VAE for latent space compression
Classical VAEs for image and video compression include standard VAEs(Kingma, 2013), quantized versions like VQVAE(Esser et al., 2021), VQGAN(Van Den Oord et al., 2017), and their GAN-enhanced variants, which improve reconstruction quality for higher compression. For example, VideoGPT(Yan et al., 2021) use a 3D-VQVAE in the latent space for video generation. Seine(Chen et al., 2024e) and MAGVIT(Yu et al., 2023a) integrate 3D-VQGAN with discriminators for latent space encoding to achieve better visual quality. The VAE framework not only compresses effectively but also enable the training of multiple generative tasks across downstream tasks.
@@
By integrating VAE in model building, some image generation models(Ho et al., 2022a; Li et al., 2024b; Podell et al., 2023; Bao et al., 2023b; Peebles & Xie, 2023; Lu et al., 2024c; Ma et al., 2024a; Chen et al., 2023c; Li et al., 2024l; Team, 2024b; Esser et al., 2024) also leverage latent space encoding/decoding, freezing VAE parameters during diffusion training and inference. Some diffusion models, including Make-A-Video(Singer et al., 2022), Imagen(Ho et al., 2022a), Show-1(Zhang et al., 2023a), directly learn pixel distributions generating videos, as shown in Table 1. However, video generation involves both spatial and temporal complexity, leading to higher computational costs. Besides, some representative diffusion video generation models, including Sora(Liu et al., 2024e), Phenaki(Villegas et al., 2022), VideoCrafter(Chen et al., 2023a), AnimateDiff(Guo et al., 2023c), and VideoPoet(Kondratyuk et al., 2023), start using VAE compression in video settings before training and inference in latent space. Many video generation models, including Latte(Ma et al., 2024c), Lavie(Huang et al., 2023b), MagicVideo(Zhou et al., 2022), Align-your-latent(Blattmann et al., 2023b), and AnimateDiff(Guo et al., 2023c), are devrived from Stable Diffusion’s image 2D VAEs, as training a full 3D latent space from scratch is challenging. However, temporal compression in 2D VAE-based video models relies on uniform frame sampling, which ignores motion information, leading to low FPS and less smooth video generation.
Some approaches also start to utilize hybrid 2D-3D or fully 3D VAEs. For example, MagViT (Yu et al., 2023a), EasyAnimate (Xu et al., 2024c), Open-Sora (Zheng et al., 2024c; Lab & etc., 2024), and CogVideo-X (Yang et al., 2024j) employ hybrid 2D-3D VAEs, leveraging volumetric encoding for video generation. MAGViT(Yu et al., 2023a) adopting 3D VQGAN structures integrating both 3D and 2D downsampling, and MAGViT-V2(Yu et al., 2024b) utilizing a fully 3D VAE with 3D convolutional encoder and overlapping downsampling for enhanced spatial-temporal fidelity. The MAGVIT architecture improves video compression by combining spatial and temporal information in one model. This helps reduce data size while keeping good video quality.
@@
HunyuanVideo(Kong et al., 2025), Kling, Mochi 1 (Team, 2024a), and SkyReels-V1(SkyReels-AI, 2025) continue leveraging 3D VAEs for compressed training in latent space in industry scenarios. Step-Video-T2V(Ma et al., 2025) and Magic-1-For-1(Yi et al., 2025) introduce novel Video VAE structures for efficient generation. However, training a video VAE independently without ensuring compatibility often leads to a latent space gap, preventing accurate projection into pixel space (Zhao et al., 2024b). CV-VAE (Zhao et al., 2024b) introduces a novel approach using latent space regularization to train a video VAE that extracts continuous latent for generative video models while maintaining compatibility with existing pretrained image and video models. Also, to trade off lower memory and computational cost with slightly lower reconstruction quality, the latest video generation model, Moive Gen(teamMeta, 2024) further uses interleaved 2D-1D convolutional encoders in its VAE. This approach balances efficiency and quality by reducing the redundancy of temporal information while maintaining spatial coherence.
Figure 7: 3D causal VAE structure in CogVideoX (Yang et al., 2024j). (a) The encoder compresses spatial and temporal features, followed by spatial-only layers and a KL regularizer to generate latent representations. The decoder mirrors this structure in reverse. (b) Temporally causal 3D convolutions are applied so each output frame depends only on the current and earlier frames, maintaining motion consistency.
As video generation increasingly demands temporal consistency and causal modeling, recent approaches have introduced causal 3D VAE structures that process frames in a strictly sequential, past-dependent manner. As shown in Figure 7, CogVideoX (Yang et al., 2024j) uses a 3D VAE with temporal causality to compress spatial and temporal information, ensuring that each output frame is based only on current and earlier frames.
In the encoder-decoder pipeline, the model applies downsampling across both space and time, then reconstructs frames with mirrored upsampling layers. The VAE operates in a causal, frame-by-frame fashion: it never uses future frames when encoding or decoding the present one, which helps ensure temporal coherence in longer video outputs.
2.5.4 Text encoders